# Spin-Peierls transition in the frustrated spinels ZnCr2O4 and MgCr2O4

### v1

@@ Abstract
In this letter, we develop an original approach to understand this transition straddling frustrated magnetism and magneto-elastic distortion, using a combination of Density Functional Theory (DFT) and Monte Carlo simulations, based on experimental data. First, we compute the ab initio spin exchange Hamiltonian of these two materials for the three distinct crystal structures that have been found in experiments: cubic ($Fd\bar{3}m$), tetragonal ($I4_{1}/amd$) and orthorhombic ($Fddd$). MgCr2O4 and ZnCr2O4 are known to be proximate of the pyrochlore antiferromagnetic spin liquid, whose extensive entropy strongly hinders any ordering mechanism. Quantitative results around the transition thus necessarily require an approach beyond mean field: We simulate these ab initio Hamiltonians with a Monte Carlo algorithm, able to properly account for the thermal fluctuations of this strongly correlated magnetic texture. Our main results are that, despite an exchange energy scale of $\theta_{\rm CW}\sim 400$ K, and distinct Hamiltonians derived from independent experimental data, our simulations remarkably fit the experimental transition temperature for both materials within a few Kelvin only. The models with cubic (resp. tetragonal and orthorhombic) symmetry have a transition temperature below (resp. above) the one found in experiments, $T_{\rm N}\approx 12.5\pm 0.5$ K. In addition, we confirm that the low-symmetry systems are magnetically more stable at $T_{\rm N}$. Our results are consistent with a spin-Peierls scenario where the cubic symmetry persists down to $T_{\rm N}$, where the structural transition into a lattice with lower symmetry is induced by a gain in magnetic exchange energy and accompanied by long-range magnetic order at the same temperature $T_{\rm N}$.
@@
We perform DFT energy mapping for the six crystal structures in order to determine the parameters of the Heisenberg Hamiltonian (1). This approach has already been successful for other Cr3+ materials with $S=3/2$ [21, 22, 23]. All electronic structure calculations are based on the full potential local orbital (FPLO) method [24] in combination with a generalized gradient approximation (GGA) exchange correlation functional [25]. We determine the exchange interactions by calculating DFT energies for a number of spin configurations that is substantially larger than the number of couplings we want to determine. This requires a $2\times 1\times 1$ supercell with eight symmetry inequivalent Cr moments in the case of the cubic structure and $2\times 2\times 1$ supercells with sixteen symmetry inequivalent Cr moments in the case of tetragonal and orthorhombic structures. The results are shown graphically in Fig. 2. In order to determine the experimentally appropriate value of the $U$ parameter for Cr $3d$ in the DFT+U calculations [26], we demand that the Hamiltonian couplings match the experimentally determined Curie-Weiss temperatures $\theta_{\rm CW}=-390$ K for ZnCr2O4 [19, 2] and $\theta_{\rm CW}=-400$ K for MgCr2O4 [5, 2]. The Curie-Weiss temperatures are estimated for the cubic structure as
@@
The selected $U$ values are between 2 and 2.5 eV and comparable with values found to be relevant for other chromium spinels [22]. There is a certain variation in the reported Curie-Weiss temperatures of ZnCr2O4 [27, 28, 19, 29] and MgCr2O4, but as the evolution of exchange interactions with $U$ is rather smooth and ratios do not vary much, the results are not very sensitive to the choice. For the cubic structure, comparable calculations have been performed before [30, 31, 16]. However, for ZnCr2O4, Ref. 30 does not separate $J_{3a}$ and $J_{3b}$, and Ref. 31 reports identical Hamiltonians for cubic ZnCr2O4 and MgCr2O4. Our results for cubic MgCr2O4 are consistent with ab-initio calculations of Ref. [16] and can be compared to the determination of the Heisenberg Hamiltonian parameters by Bai et al[13] using fits to inelastic neutron scattering data; they find the set of couplings $\{J_{1},J_{2},J_{3a},J_{3b}\}$ to be $\{38.05(3),3.10,4.00,0.32\}$ K. We do not expect to find a perfect match with ab-initio calculations, but it is particularly encouraging for our DFT results to recover the same main physics: We also find a dominant antiferromagnetic $J_{1}$ coupling, one order of magnitude larger than the second exchange path $J_{3a}$ (also antiferromagnetic) and with a small $J_{3b}$ term. There is a difference though, in that the $J_{2}$ term changes sign. This exchange path is, however, known to be subdominant, with previous studies disagreeing on its sign [31, 16, 13]. For the sake of completeness, we have checked that using Bai’s parameters does not affect the conclusions of our paper [32]. In addition the structure factor obtained by Monte Carlo simulations at 20 K [see discussion below and Fig. 3(e)] is consistent with the one measured experimentally in [13]. We are thus confident that our ab-initio results are set on a sensible experimental footing. To the best of our knowledge, no ab-initio estimates of the tetragonal and orthorhombic Hamiltonian couplings have been published for ZnCr2O4 and MgCr2O4. We chose to use the same value of $\theta_{\rm CW}$ for the three lattice structures of each material as it allows us to impose a constant energy scale of the total magnetic exchange path across the transition. Our results are thus a direct consequence of the lower degree of magnetic frustration in lower-symmetry structures, and cannot be due to an overall rescaling of the exchange couplings.
@@
The transition temperatures for cubic simulations are lower than the one found in experiments $T_{\rm N}=12.5\pm 0.5$ K. Simulations thus expect ZnCr2O4 and MgCr2O4 to remain magnetically disordered down to $T_{\rm N}$, as observed in experiments. On the other hand, the transition temperatures for tetragonal and orthorhombic simulations are higher than $T_{\rm N}$. It means that a structural transition in this temperature range is automatically accompanied by magnetic ordering, as observed in experiments. And at $T_{\rm N}$, the magnetic energy of the system is always lower for the tetragonal and orthorhombic structures than for the cubic one. This difference in magnetic energy provides an approximate estimate of the electrostatic energy cost due to lattice distortions.
Now, keeping in mind that we are simulating systems whose exchange energy scales are of the order of $|\theta_{CW}|\approx 400$ K, and whose coupling parameters have been obtained from unbiased DFT calculations derived from independent experimental data (X-ray diffraction and magnetic susceptibility measurements) for two different materials, it is truly remarkable that the transition temperatures of the cubic and tetragonal structures occur within a very small temperature window $\Delta T\sim 5{\rm K}\ll|\theta_{CW}|$, with the experimental transition temperature $T_{\rm N}$ sitting precisely between them. To this day, the literature remains unclear whether both tetragonal and orthorhombic structures really co-exist in ZnCr2O4 and MgCr2O4, and in which ratio. Our simulations suggest that the tetragonal distortion might play a key role in the spin-Peierls mechanism. Indeed, both ZnCr2O4 and MgCr2O4 essentially remain magnetically disordered down to the necessary temperature for the tetragonal distortion to be able to bring a net magnetic energy gain [Fig. 3(c,d)] that nucleates the structural transition. In that sense, the presence of an orthorhombic distortion in these materials might be mediated by the tetragonal one.
Another interesting outcome is that the cubic magnetic transition is only a few Kelvin below $T_{\rm N}$, i.e., only a few Kelvin below the experimentally accessible regime with cubic structure. It is thus not unreasonable for Ref. [34] to have recently observed an onset of magnetic order above $T_{N}$ in MgCr2O4.
@@
Acknowledgements–The work of Y.I. was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-2210452. The participation of Y.I. at the Aspen Center for Physics was supported by the Simons Foundation. The research of Y.I. was carried out, in part, at the Kavli Institute for Theoretical Physics in Santa Barbara during the “A New Spin on Quantum Magnets” program in summer 2023 and “Correlated Gapless Quantum Matter” program in spring 2024, supported by the National Science Foundation under Grant No. NSF PHY-1748958. Y.I. acknowledges support from the ICTP through the Associates Programme, from the Simons Foundation through Grant No. 284558FY19, and IIT Madras through the Institute of Eminence (IoE) program for establishing QuCenDiEM (Project No. SP22231244CPETWOQCDHOC). Y.I. also acknowledges the use of the computing resources at HPCE, IIT Madras. L.J. and H.O.J. thank IIT Madras for a Visiting Faculty Fellow position under the IoE program during which this work was initiated. L.J. acknowledges financial support from Grants No. ANR-18-CE30-0011-01 and ANR-23-CE30-0038-01. Y.I. and H.O.J. acknowledge the hospitality of the University of Bordeaux/CNRS during a work visit in September 2023. All authors thank the International Centre for Theoretical Sciences (ICTS), Bengaluru, India for hospitality during the program “Frustrated Metals and Insulators” (Code No. ICTS/frumi2022/9).

### v2

@@ Abstract
In this letter, we develop an original approach to understand this transition straddling frustrated magnetism and magneto-elastic distortion, using a combination of Density Functional Theory (DFT) and Monte Carlo simulations, based on experimental data. First, we compute the ab initio spin exchange Hamiltonian of these two materials for the three distinct crystal structures that have been found in experiments: cubic ($Fd\bar{3}m$), tetragonal ($I4_{1}/amd$) and orthorhombic ($Fddd$). MgCr2O4 and ZnCr2O4 are known to be close to the pyrochlore antiferromagnetic spin liquid, whose extensive entropy strongly hinders any ordering mechanism. Quantitative results around the transition thus necessarily require an approach beyond mean field: We simulate these ab initio Hamiltonians with a Monte Carlo algorithm, able to properly account for the thermal fluctuations of this strongly correlated magnetic texture. Our main results are that, despite an exchange energy scale of $\theta_{\rm CW}\sim 400$ K, and distinct Hamiltonians derived from independent experimental data, our simulations remarkably fit the experimental transition temperature for both materials within a few Kelvin only. The models with cubic (resp. tetragonal and orthorhombic) symmetry have a transition temperature below (resp. above) the one found in experiments, $T_{\rm N}\approx 12.5\pm 0.5$ K. In addition, we confirm that the low-symmetry systems are magnetically more stable at $T_{\rm N}$. Our results are consistent with a spin-Peierls scenario where the cubic symmetry persists down to $T_{\rm N}$, where the structural transition into a lattice with lower symmetry is induced by a gain in magnetic exchange energy and accompanied by long-range magnetic order at the same temperature $T_{\rm N}$.
@@
We perform DFT energy mapping for the six crystal structures in order to determine the parameters of the Heisenberg Hamiltonian (1). This approach has already been successful for other Cr3+ materials with $S=3/2$ [21, 22, 23]. All electronic structure calculations are based on the full potential local orbital (FPLO) method [24] in combination with a generalized gradient approximation (GGA) exchange correlation functional [25]. We use collinear DFT here; effects of spin-orbit coupling like the Dzyaloshinskii-Moriya (DM) interaction that would require non-collinear DFT can be expected to be small in Cr3+ due to the half-filled $t_{2g}$ shell, and a calculation for MgCr2O4 has found DM to be negligible [16], while previous theoretical fits to neutron-scattering measurements did not require any anisotropic terms [26, 13]. At lower energy scale, anisotropy could possibly play a role to understand the subtle degeneracy lifting [13, 27] responsible for the magnetic textures at very low temperatures. While this question is beyond the scope of our present Letter, non-collinear DFT methods [28] could be useful in that context. We determine the exchange interactions by calculating DFT energies for a number of spin configurations that is substantially larger than the number of couplings we want to determine. This requires a $2\times 1\times 1$ supercell with eight symmetry inequivalent Cr moments in the case of the cubic structure and $2\times 2\times 1$ supercells with sixteen symmetry inequivalent Cr moments in the case of tetragonal and orthorhombic structures. The results are shown graphically in Fig. 2. In order to determine the experimentally appropriate value of the $U$ parameter for Cr $3d$ in the DFT+U calculations [29], we demand that the Hamiltonian couplings match the experimentally determined Curie-Weiss temperatures $\theta_{\rm CW}=-390$ K for ZnCr2O4 [19, 2] and $\theta_{\rm CW}=-400$ K for MgCr2O4 [5, 2]. The Curie-Weiss temperatures are estimated for the cubic structure as
@@
The selected $U$ values are between 2 and 2.5 eV and comparable with values found to be relevant for other chromium spinels [22]. There is a certain variation in the reported Curie-Weiss temperatures of ZnCr2O4 [30, 31, 19, 32] and MgCr2O4, but as the evolution of exchange interactions with $U$ is rather smooth and ratios do not vary much, the results are not very sensitive to the choice. For the cubic structure, comparable calculations have been performed before [33, 34, 16]. However, for ZnCr2O4, Ref. 33 does not separate $J_{3a}$ and $J_{3b}$, and Ref. 34 reports identical Hamiltonians for cubic ZnCr2O4 and MgCr2O4. Our results for cubic MgCr2O4 are consistent with ab-initio calculations of Ref. [16] and can be compared to the determination of the Heisenberg Hamiltonian parameters by Bai et al[13] using fits to inelastic neutron scattering data; they find the set of couplings $\{J_{1},J_{2},J_{3a},J_{3b}\}$ to be $\{38.05(3),3.10,4.00,0.32\}$ K. We do not expect to find a perfect match with ab-initio calculations, but it is particularly encouraging for our DFT results to recover the same main physics: We also find a dominant antiferromagnetic $J_{1}$ coupling, one order of magnitude larger than the second exchange path $J_{3a}$ (also antiferromagnetic) and with a small $J_{3b}$ term. There is a difference though, in that the $J_{2}$ term changes sign. This exchange path is, however, known to be subdominant, with previous studies disagreeing on its sign [34, 16, 13]. For the sake of completeness, we have checked that using Bai’s parameters does not affect the conclusions of our paper [35]. In addition the structure factor obtained by Monte Carlo simulations at 20 K [see discussion below and Fig. 3(e)] is consistent with the one measured experimentally in [13]. We are thus confident that our ab-initio results are set on a sensible experimental footing. To the best of our knowledge, no ab-initio estimates of the tetragonal and orthorhombic Hamiltonian couplings have been published for ZnCr2O4 and MgCr2O4. We chose to use the same value of $\theta_{\rm CW}$ for the three lattice structures of each material as it allows us to impose a constant energy scale of the total magnetic exchange path across the transition. Our results are thus a direct consequence of the lower degree of magnetic frustration in lower-symmetry structures, and cannot be due to an overall rescaling of the exchange couplings.
@@
The transition temperatures for cubic simulations are lower than the one found in experiments $T_{\rm N}=12.5\pm 0.5$ K. Simulations thus expect ZnCr2O4 and MgCr2O4 to remain magnetically disordered down to $T_{\rm N}$, as observed in experiments. On the other hand, the transition temperatures for tetragonal and orthorhombic simulations are higher than $T_{\rm N}$. It means that a structural transition in this temperature range is automatically accompanied by magnetic ordering, as observed in experiments. And at $T_{\rm N}$, the magnetic energy of the system is always lower for the tetragonal and orthorhombic structures than for the cubic one.
In frustrated magnets, the value of $T_{\rm N}$ sensitively depends on the ratios of the exchange interactions, and is largely independent from $\theta_{\rm CW}$; this is especially true for ZnCr2O4 and MgCr2O4 which are very close to the Heisenberg antiferromagnetic spin liquid. In that context, and keeping in mind that we are simulating systems whose exchange energy scales are of the order of $|\theta_{CW}|\approx 400$ K, and whose coupling parameters have been obtained from unbiased DFT calculations derived from independent experimental data (X-ray diffraction and magnetic susceptibility measurements) for two different materials, it is truly remarkable that the transition temperatures of the cubic and tetragonal structures occur within a very small temperature window $\Delta T\sim 5{\rm K}\ll|\theta_{CW}|$, with the experimental transition temperature $T_{\rm N}$ sitting precisely between them. Our simulations show that both ZnCr2O4 and MgCr2O4 remain magnetically disordered down to the necessary temperature for the tetragonal distortion to be able to bring a net magnetic energy gain [Fig. 3(c,d)]. This is consistent with a three-dimensional analog of the spin-Peierls mechanism, where the tetragonal distortion is made possible thanks to the energy gain in magnetic interactions. To this day, the literature remains unclear whether both tetragonal and orthorhombic structures really co-exist, and in which ratio and temperature window. But since the simulated transition temperatures of the orthorhombic structures, respectively 70 K and 25 K for ZnCr2O4 and MgCr2O4, are noticeably higher than in experiments, our simulations suggest that the spin-Peierls mechanism is attached to the tetragonal distortion. In the future, it would be interesting to see if the tetragonal structure could serve as an intermediate state facilitating the onset of an orthorhombic distortion.
Another notable outcome is that the cubic magnetic transition is only a few Kelvin below $T_{\rm N}$, i.e., only a few Kelvin below the experimentally accessible regime with cubic structure. It is thus not unreasonable for Ref. [36] to have recently observed an onset of magnetic order above $T_{N}$ in MgCr2O4.
@@
Acknowledgements–The work of Y.I. was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-2210452. The participation of Y.I. at the Aspen Center for Physics was supported by the Simons Foundation (1161654, Troyer). The research of Y.I. was carried out, in part, at the Kavli Institute for Theoretical Physics in Santa Barbara during the “A New Spin on Quantum Magnets” program in summer 2023 and “Correlated Gapless Quantum Matter” program in spring 2024, supported by the National Science Foundation under Grant No. NSF PHY-2309135. Y.I. acknowledges support from the ICTP through the Associates Programme, from the Simons Foundation through Grant No. 284558FY19, and IIT Madras through the Institute of Eminence (IoE) program for establishing QuCenDiEM (Project No. SP22231244CPETWOQCDHOC). Y.I. also acknowledges the use of the computing resources at HPCE, IIT Madras. L.J. and H.O.J. thank IIT Madras for a Visiting Faculty Fellow position under the IoE program during which this work was initiated. L.J. acknowledges financial support from Grants No. ANR-18-CE30-0011-01 and ANR-23-CE30-0038-01. Y.I. and H.O.J. acknowledge the hospitality of the University of Bordeaux/CNRS during a work visit in September 2023. All authors thank the International Centre for Theoretical Sciences (ICTS), Bengaluru, India for hospitality during the program “Frustrated Metals and Insulators” (Code No. ICTS/frumi2022/9).